### v1

# From minimal-length quantum theory to modified gravity

Abstract
In this work, we consider generalized uncertainty principles (GUPs) that incorporate a minimal length through generic momentum-dependent deformation functions. We thus develop a systematic approach connecting such a framework to effective gravitational actions extending general relativity. By examining quantum gravity-motivated corrections to black hole entropy induced by the GUP and employing Wald’s formalism, we reconstruct modifications to Einstein’s gravity within the contexts of $f(R)$ and $f(R,R_{\mu\nu}R^{\mu\nu})$ theories. In this way, we establish a direct mapping between the GUP parameters and the higher-order curvature coefficients in the gravitational Lagrangian. As an illustrative application, we compute corrections to the general relativistic prediction for light deflection, which in turn allows us to infer a stringent upper bound on the minimal measurable length. Our results show that GUP-induced effects can be consistently embedded into extended gravity theories, offering a promising framework for testing quantum gravity phenomenology through astrophysical and cosmological observations.
I Introduction
The search for observable imprints of quantum gravity in semi-classical regimes has led to the development of various phenomenological models. Among them, the generalized uncertainty principle (GUP) has emerged as a prominent candidate for encoding Planck-scale corrections to quantum mechanics. By modifying the standard Heisenberg uncertainty relation, the GUP introduces a momentum-dependent deformation that incorporates a minimal measurable length [Kempf:1994su, Nozari:2005mr, Das:2009hs, Pedram:2011xj, Hossenfelder:2012jw, Bosso:2023sxr]. This minimal scale, often associated with the Planck length, is a common feature in several approaches to quantum gravity, including string theory, loop quantum gravity and non-commutative geometry [Gross:Mende:1988, AmelinoCamelia:2002dsr, Rovelli:1998lrr, Maggiore:1993rv, Konishi:1989wk, Capozziello:1999wx, Bosso:2024nmn].
On the gravitational side, classical general relativity (GR) can be generalized by including higher-order curvature invariants in the action. Such extensions give rise to theories such as $f(R)$ gravity [Starobinsky:1980te, Barrow:1988xh, Capozziello:2003tk, Clifton:2005aj, Starobinsky:2007hu, Sotiriou:2008rp, DeFelice:2010aj, Nojiri:2017ncd, Capozziello:2019cav, Bajardi:2022ocw], or more broadly to Lagrangians depending on both the Ricci scalar and tensor invariants, as in $f(R,R_{\mu\nu}R^{\mu\nu})$ gravity [Carroll:2004de, Bogdanos:2009tn, Bajardi:2023shq]. These modifications are not introduced arbitrarily; rather, they emerge naturally within effective field theory treatments of gravity coupled to quantum matter fields [Birrell:1982ix], and they are supported by the renormalization structure of quantum field theory in curved spacetime [Sotiriou:2008rp, Nojiri:2010wj].
In the absence of a complete and experimentally confirmed theory of quantum gravity, one viable strategy is to rely on phenomenological approaches that can bridge quantum mechanical modifications with semi-classical gravitational dynamics. In this perspective, models such as the GUP provide a useful parametrization of quantum gravitational effects at low energies, suggesting how Planck-scale features might imprint themselves in observable regimes.
Several studies have investigated how GUP-induced corrections affect black hole (BH) physics [Scardigli:1999jh, Buoninfante:2019fwr, Adler:2001vs, Medved:2004yu, Custodio:2003jp], with particular focus on the entropy–area relation [Majumder:2011wl, Bosso:2023fnb, Buoninfante:2020guu, Tawfik:2015kga]. Such corrections typically involve logarithmic and inverse-area terms, which are likewise predicted by higher-curvature gravity theories. This formal resemblance suggests a deeper structural connection between the two frameworks and motivates a more systematic investigation of their correspondence. Despite these advances, a general methodology for reconstructing effective gravitational Lagrangians from arbitrary GUP deformations remains lacking. Most existing analyses, in fact, focus on particular GUP forms and a limited class of gravitational models [Hammad:2015dka], often without providing a unified procedure that connects the deformation function to curvature corrections at the action level.
In this paper, we propose a general framework to establish this connection. Starting from a generic expansion of the GUP deformation function in powers of momentum uncertainty, we derive the corresponding corrections to BH entropy by interpreting the GUP-modified uncertainty as a regulator near the horizon. These entropy corrections are then translated into modifications of the gravitational action using Wald’s entropy formula, applied first to $f(R)$ and then to the more general $f(R,R_{\mu\nu}R^{\mu\nu})$ class of theories. By inverting the resulting relations, we obtain an explicit reconstruction of the effective Lagrangian, identifying the curvature-dependent terms that encode the underlying quantum uncertainty structure. As such, our approach provides a clear and physically motivated dictionary connecting quantum mechanical deformations to semi-classical gravitational dynamics. Furthermore, it represents a natural and direct extension of the results of [Hammad:2015dka] to the fully general case of polynomial GUP models with deformation functions of arbitrary degree.
Importantly, our framework is not merely a formal or speculative correspondence: it leads to concrete physical predictions. We explicitly carry the GUP parameters into the gravitational sector and derive their phenomenological implications in the weak-field regime. In particular, the analysis of light deflection by compact objects yields novel empirical bounds on the minimal length scale, tightening and complementing existing constraints in the literature.
The structure of the paper is as follows. In Sec. II, we introduce our GUP model and examine its implications for BH entropy. In Sec. III, employing Wald’s formalism, we reconstruct the effective gravitational action that reproduces the GUP-corrected entropy within the frameworks of $f(R)$ and $f(R,R_{\mu\nu}R^{\mu\nu})$ gravity. Section IV establishes an explicit correspondence between the GUP parameters and the coefficients in the curvature expansion of the gravitational Lagrangian, while Sec. V is devoted to astrophysical applications of the obtained gravity theory. Finally, in Sec. VI, we present our conclusions and outline future directions.
Throughout this work, we adopt natural units, $k_{B}=\hbar=c=1$, such that the Planck length is expressed as $\ell_{P}=\sqrt{G}$, where $G$ denotes Newton’s gravitational constant.
II Generalized Uncertainty Principle and Black hole Entropy
The GUP is a phenomenological modification of the standard Heisenberg uncertainty relation that incorporates effects expected from quantum gravity theories. It generally introduces a minimal measurable length scale, typically on the order of the Planck length, which modifies the standard commutation relations between position and momentum operators [Kempf:1994su].
A commonly studied class of models introduces deformations that depend on powers of the momentum uncertainty or, alternatively, on the momentum operator, with the two formulations related through the Robertson-Schrödinger relation. A well-known example is the model proposed in [Kempf:1994su], in which the canonical commutator is deformed as $x,p]=i(1+\beta p^{2})$, where $\beta$ is a small deformation parameter, typically associated with the (squared) Planck scale. This modification leads to an uncertainty relation of the form $\Delta x\Delta p\geq\frac{1}{2}\left[1+\beta(\Delta p)^{2}\right]$, which implies the existence of a non-zero minimal uncertainty in position, $\Delta x_{\text{min}}=\sqrt{\beta}>0$. The physical consequences of this feature have been widely discussed in the literature [Scardigli:2003kr, Jizba:2009qf, Amelino-Camelia:2005zpp, Das:2008kaa, Carr:2015nqa, Scardigli:2018jlm, Anacleto:2015mma, Jizba:2022icu, Tawfik:2015rva, Barca:2023shv, Segreto:2024gcg] (see also [Bosso:2023aht] for a recent review).
Alternative formulations based on deformed commutators have also been proposed, such as $x,p]=i\,e^{\beta p^{2}} [$Nouicer:2007jg] or $x,p]=i(1-\beta p^{2})^{-1} [$Pedram:2011gw], which represent non-polynomial generalizations of the canonical commutation relation and lead to qualitatively different physical implications. The use of even powers, such as $p^{2}$ or $(\Delta p)^{2}$, is motivated primarily by symmetry considerations. In particular, these functions are invariant under parity transformations $p\rightarrow-p$, preserving the isotropy of space and ensuring that the deformation respects fundamental discrete symmetries. However, more exotic models have been put forward in which the deformation also includes odd powers of the momentum, such as a linear dependence [Ali:2011fa]. These terms, which explicitly break parity symmetry, may emerge in theories with intrinsic chirality, anisotropic backgrounds or non-commutative geometries.
To keep the analysis as general as possible, we adopt a framework in which the deformation is represented by an unspecified analytic function $\Xi(\Delta p)$, which admits a power series expansion, namely
$\Delta x\,\Delta p\geq\frac{1}{2}\left[1+\Xi(\Delta p)\right],$ (1)
with
$\Xi(\Delta p)=\sum_{k=1}^{\infty}{\ell_{P}^{k}}c_{k}(\Delta p)^{k},$ (2)
where $c_{k}$ are dimensionless coefficients and $\ell_{P}$ denotes the Planck length. This form ensures that higher-order contributions are suppressed by powers of the Planck scale, so that standard quantum mechanics is recovered in the low-energy regime $\Delta p\ll\ell_{P}^{-1}$. It is straightforward to verify that, in the limit where all coefficients $c_{k}$ vanish except for $c_{2}$, the above model reduces to the quadratic GUP commonly employed in the literature.
In order to proceed with analytical calculations, it is customary to assume that inequality (1) is approximately saturated. This approximation corresponds to the coherent states of the ordinary theory, i.e., states characterized by a minimal uncertainty product. Such states, in the framework of minimal-length quantum mechanics, have been extensively considered in the literature (see, e.g., [Kempf:1994su, Pedram:2012ui, Dey:2012tv, Bosso:2017ndq, Bosso:2021koi, Jizba:2023ygi]). From a computational perspective, the use of an approximate equality is essential to derive closed-form expressions or to make progress in perturbative expansions. Moreover, many physical applications of the GUP, such as in BH thermodynamics, quantum cosmology or minimal length models, rely on this approximation to extract meaningful phenomenological consequences [Hossenfelder:2012jw].
Under the above assumption, from the GUP (1) one obtains an implicit relation for $\Delta p$ in terms of $\Delta x$:
$\Delta p=\frac{1}{2\Delta x}\left[1+\Xi(\Delta p)\right].$ (3)
Inverting this relation yields
$\Delta p=\sum_{n=0}^{\infty}\ell_{P}^{n}\frac{\tilde{c}_{n}}{(2\Delta x)^{n+1}}\,$, (4)
with $\tilde{c}_{0}=1$ and
$\tilde{c}_{n}=\sum_{\vec{m}\in P(n)}\frac{n^{\underline{|\vec{m}|-1}}}{\vec{m}!}\vec{c}^{\vec{m}}\,,\quad n>0\,.$ (5)
The definitions of all symbols appearing in the above relation, along with the computational details of its derivation, are provided in Appendix A. By way of illustration, we report below the expressions for the first five coefficients:
$\tilde{c}_{1}$ $=c_{1}\,$, (6a)
$\tilde{c}_{2}$ $=c_{1}^{2}+c_{2}\,$, (6b)
$\tilde{c}_{3}$ $=c_{1}^{3}+3c_{1}c_{2}+c_{3}\,$, (6c)
$\tilde{c}_{4}$ $=c_{1}^{4}+6c_{1}^{2}c_{2}+2c_{2}^{2}+4c_{1}c_{3}+c_{4}\,$, (6d)
$\tilde{c}_{5}$ $=c_{1}^{5}+10c_{1}^{3}c_{2}+10c_{1}c_{2}^{2}+10c_{1}^{2}c_{3}+5c_{2}c_{3}+5c_{1}c_{4}+c_{5}\,.$ (6e)
II.1 Implications for BH entropy
As discussed earlier, the GUP modifies the standard Heisenberg relation by introducing a minimal measurable length, typically of the order of the Planck scale. Such a feature is expected in any consistent theory of quantum gravity, as it prevents the probing of distances below the Planck scale where classical concepts of spacetime break down. BH horizons represent the natural arena to investigate these effects, since the relevant phenomena occurring in their vicinity (such as Hawking radiation) are governed by quantum gravity laws. In particular, the typical size of quantum fluctuations of matter fields cannot be smaller than this minimal length scale. Therefore, the absorption or emission of a particle by a BH is necessarily constrained by this fundamental limit.
This constraint directly affects the properties of the BH horizon area. When a particle is absorbed, the minimal increase in the horizon area $\Delta A$ is linked to the particle’s position uncertainty $\Delta x$ [Awad:2014bta, Hammad:2015dka]. Incorporating the GUP-modified uncertainty leads to corrections in the quantization of horizon area and consequently to corrections in the Bekenstein-Hawking entropy formula:
$S=\frac{A}{4G}+S_{\mathrm{corr}}\,$, (7)
where $S_{\mathrm{corr}}$ contains sub-leading terms due to quantum gravity effects. On the other hand, starting from BH absorption and emission, with a Gedankenexperiment similar to Heisenberg’s microscope, one can arrive at a GUP formulation [Maggiore:1993rv].
As stated above, the heuristic derivation of $S_{\mathrm{corr}}$ proceeds by estimating the minimal area increase $\Delta A$ caused by absorption of a quantum particle with position uncertainty $\Delta x$ and momentum uncertainty $\Delta p$. Since the energy of a particle with size $\Delta x$ near the horizon can be estimated by $E\sim\Delta p$ and considering that the minimal increase in BH area associated with the absorption of such a particle is $\Delta A_{\text{min}}\sim GE\Delta x$ [Hammad:2015dka], we obtain from Eq. (4)
$\Delta A_{\min}=\lambda G\left[1+\sum_{k=1}G^{k/2}\tilde{c}_{k}(2\Delta x)^{-k}\right],$ (8)
where we have used $\ell_{P}^{2}=G$ and all proportionality factors have been absorbed into the coefficient $\lambda$, which is to be determined. Hereafter, unless otherwise specified, the summation index $k$ runs from $1$ to $\infty$.
Replacing $\Delta x=\zeta\sqrt{A}$, where $\zeta$ is a dimensionless proportionality constant, Eq. (8) can be expressed as a function of the horizon area:
$\Delta A_{\min}=\lambda G\left(1+\sum_{k}d_{k}A^{-k/2}\right),$ (9)
where the coefficients $d_{k}$ are directly related to the deformation parameters $c_{k}$ through $d_{k}=G^{k/2}\left(2\zeta\right)^{-k}\tilde{c_{k}}$.
The entropy is then interpreted as counting the number of fundamental area units $\Delta A_{\min}$ that compose the total BH horizon area:
$S_{\text{GUP}}=\int\frac{dA}{\Delta A_{\min}}=\frac{1}{\lambda G}\int\frac{dA}{1+\sum_{k}d_{k}A^{-k/2}}.$ (10)
Expanding for $\sum_{k}d_{k}A^{-k/2}\ll 1$ yields
$S_{\text{GUP}}=\frac{1}{\lambda G}\int dA\left(1-\sum_{k}d_{k}A^{-k/2}\right),$ (11)
and integrating term by term, we have
$S_{\text{GUP}}=\frac{1}{\lambda G}\left(A-2d_{1}\sqrt{A}-d_{2}\ln A+\sum_{k\geq 3}\frac{2d_{k}}{k-2}A^{1-k/2}\right),$ (12)
up to an integration constant, which can be set to restore the correct dimensions of the argument of the logarithm. We notice that, regardless of the sign of the coefficients $d_{k}$, the modified entropy in Eq. (12) is expected to remain positive, since the corrections induced by the GUP are subdominant compared to the leading-order term $A/(4G)$ (see the discussion below Eq. (2)).
Moreover, we see that we should have $\lambda=4$ in order to recover the standard Bekenstein-Hawking entropy in the limit where $d_{k}=0$, $\forall k$. Thus, the final GUP entropy formula takes the form:
$S_{\text{GUP}}=\frac{A}{4G}+a_{1}\sqrt{A}-a_{2}\ln A+\sum_{k\geq 3}a_{k}\,A^{1-k/2}\,$, (13)
where, to simplify the notation, we have defined
$a_{k}=\begin{dcases}\dfrac{d_{2}}{4G},&\text{if }k=2\,,\\[5.69054pt] \dfrac{d_{k}}{2G(k-2)},&\text{if }k\neq 2\,.\end{dcases}$ (14)
Equation (13) shows the generic structure of the entropy once GUP corrections are taken into account. A crucial point is that, in general, there is no one-to-one mapping between the coefficients $a_{k}$ in $S_{\text{GUP}}$ and the coefficients $c_{k}$ in Eq. (1). For instance, the logarithmic correction connected to $a_{2}$ can originate from both linear and quadratic terms in the GUP (cf. Eq. (6b)), and analogous considerations hold for higher-order corrections.
Nevertheless, if for each $k$ one considers only the leading corrections [^1] For example, if only $c_{1}\neq 0$, the leading-order term is given by Eq. (6a), as $c_{1}^{n}$ with $n>1$ are subdominant and, thus, $\tilde{c}_{k}=0,\,\forall k>1$., one may infer a direct correspondence between the terms in the GUP expansion and the corrections to the entropy. For instance, we observe that if $c_{k}=0,\,\forall k\neq 1$, the Bekenstein-Hawking entropy acquires a correction that scales as $\sqrt{A}$, which is indicative of a genuinely infrared (IR) effect. This outcome is consistent with results obtained in other approaches involving linear GUP deformations; see, for example, [Majumder:2012rtc, Tawfik:2015kga]. Furthermore, it has been identified in non-commutative quantization contexts [Nozari:2008rc]. It is therefore evident that, when focusing on the ultraviolet (UV) regime of quantum gravity, such a term becomes subdominant compared to the leading corrections and can be safely neglected. From the GUP perspective, this provides a further argument in support of models in which the linear term is absent and the leading-order deformation is quadratic in $\Delta p$ [Kempf:1994su], as these models are more naturally associated with the existence of a minimal length and are better suited to capturing the dominant features of the UV regime.
On the other hand, if only $c_{2}\neq 0$, we have a logarithmic modification of the entropy-area relation. This term emerges as a nearly universal feature of UV quantum gravity, although different approaches yield variations in the numerical coefficient that multiplies the logarithmic contribution [Carlip:2000nv, Kaul:2000kf, Sen:2012dw, Adler:2001vs, Medved:2004yu]. Additionally, this type of correction is supported by entanglement entropy calculations in BH spacetimes in the UV regime [Solodukhin:2011gn, Alonso-Serrano:2022qvo]. Finally, for each $c_{k}\neq 0$ with $k\geq 3$, the entropy corrections scale as $A^{1-k/2}$, indicating that they become relevant only in the deep UV regime.
Overall, the corrections appearing in Eq. (12) imply that the BH entropy deviates from the area-law, reflecting minimal length effects predicted by quantum gravity. The explicit form of the coefficients $a_{k}$ provides a bridge to the microscopic parameters of the underlying quantum theory, encoded here phenomenologically in the $c_{k}$ coefficients of the GUP. Hence, by controlling the GUP deformation function $\Xi(\Delta p)$, one controls the corrections to BH entropy and, as we will see, the corresponding modifications to the gravitational action.
In the next sections, we exploit our framework to reconstruct effective gravitational Lagrangians compatible with these entropy corrections, providing a semi-classical realization of quantum gravity effects.
III Modified Gravity and Wald Entropy
Modified theories of gravity, in particular those described by a generalized action of the form
$\mathcal{S}_{g}=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\,f(R)\,$, (15)
represent a natural extension of GR. Here, $g$ is the determinant of the metric tensor, and the Lagrangian density $f(R)$ is a non-linear function of the Ricci scalar $R$, encoding higher-order curvature corrections expected from quantum gravitational or effective field theory considerations [Nojiri:2010wj, Sotiriou:2008rp].
Such extensions emerge naturally within various theoretical frameworks. For instance, quantum field theory formulated on curved spacetime induces one-loop quantum corrections that generate terms quadratic and of higher order in curvature invariants, as extensively discussed in [Birrell:1982ix]. Similarly, effective actions derived from string theory include $\alpha$-corrections, which are systematically expressed as powers of curvature tensors and represent higher-derivative modifications to classical gravity [Zwiebach:1985uq]. Additionally, phenomenological approaches aiming to explain cosmological phenomena such as dark energy and inflation often invoke modified gravitational actions featuring curvature-dependent terms [DeFelice:2010aj]. These higher-derivative corrections modify the gravitational dynamics at small (high curvature) scales and may encode signatures of the underlying quantum theory of gravity.
The Bekenstein-Hawking entropy formula is derived within the framework of classical GR. In generalized gravity theories, this relation does not hold straightforwardly due to additional curvature-dependent terms in the action. Wald formulated a general expression for BH entropy applicable to any diffeomorphism-invariant theory of gravity [Wald:1993nt]. For the class of $f(R)$ theories, the Wald entropy is given by [Iyer:1994ys]
$S_{\text{Wald}}=\frac{1}{4G}\int_{\mathcal{H}}d^{2}x\sqrt{h}\,f_{R}(R),$ (16)
where $\mathcal{H}$ denotes the BH horizon cross-section, $h$ is the induced metric on $\mathcal{H}$, and
$f_{R}(R)\equiv\frac{df(R)}{dR}.$ (17)
This formula can be understood as a modification of the gravitational coupling constant by a scalar function $f_{R}(R)$ that depends on the local curvature at the horizon.
In spherically symmetric spacetimes, the Ricci scalar evaluated at the BH horizon $R_{H}$ typically scales inversely with the square of the horizon radius, $r_{H}$:
$R_{H}=\frac{\gamma}{r_{H}^{2}}=\frac{4\pi\gamma}{A},$ (18)
where $\gamma$ is a suitable constant that we set as $\gamma=1/(4\pi)$ for simplicity. Relationship (18) allows us to express curvature corrections as functions of the horizon area $A$.
Thus, by focusing on constant curvature solutions, the Wald entropy in $f(R)$ gravity can be expressed as [DAgostino:2024sgm]
$S_{\text{Wald}}=\frac{A}{4G}f_{R}(R_{H})\,$, (19)
where we have used
$A=\int_{\mathcal{H}}d^{2}x\,\sqrt{h}=4\pi r_{H}^{2}.$ (20)
Using Eq. (18), and matching Eq. (19) with the GUP-corrected entropy expression (13), we find
$f_{R}(R)=1+4G\left[a_{1}\sqrt{R}+a_{2}R\ln R+\sum_{k\geq 3}a_{k}R^{k/2}\right].$ (21)
Integrating the last expression, we obtain the corresponding Lagrangian:
$f(R)$ $=R+4G\bigg[\frac{2a_{1}}{3}R^{3/2}+\frac{a_{2}R^{2}}{4}(\ln R^{2}-1)$
$\hskip 51.21504pt+\sum_{k\geq 3}\frac{2a_{k}}{k+2}R^{1+k/2}\bigg]+C\,$, (22)
where $C$ is an integration constant that may be interpreted as an effective cosmological constant.
This expression makes explicit how GUP-induced entropy corrections translate into non-trivial curvature terms in the gravitational action. Thus, the quantum mechanical deformation encoded by $\Xi(\Delta p)$ corresponds semi-classically to a modified gravitational Lagrangian. It is worth reminding that the consistency of this picture requires the positivity of $f_{R}(R)$ to ensure a well-defined effective gravitational coupling and avoid ghost instabilities [Sotiriou:2008rp].
III.1 Extension to $f(R,R_{\mu\nu}R^{\mu\nu})$ gravity
While the $f(R)$ gravity framework captures a significant subset of higher-curvature corrections, quantum gravity and effective field theory considerations often necessitate a more general formulation that includes additional curvature invariants, most notably the Ricci tensor squared term $R_{\mu\nu}R^{\mu\nu}$. This leads to an extended action of the form
$\mathcal{S}_{g}=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\,f(R,P)\,$, (23)
where we define $P\equiv R_{\mu\nu}R^{\mu\nu}$. The inclusion of such terms is strongly motivated by various independent lines of reasoning. From a quantum field theoretic perspective, one-loop effective actions of matter fields propagating in curved backgrounds inevitably generate curvature-squared contributions, including both $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$, with coefficients fixed by the matter content and regularization scheme employed [Birrell:1982ix, Parker:2009uva]. These terms also arise naturally in the context of renormalization group flows of gravitational couplings, where they play a central role in the construction of asymptotically safe theories of gravity [Reuter:1996cp].
Furthermore, from the viewpoint of higher-derivative gravity, the addition of the $R_{\mu\nu}R^{\mu\nu}$ invariant alters the spectrum of gravitational excitations, introducing additional degrees of freedom that can improve the ultraviolet behavior of the theory and potentially render it perturbatively renormalizable [Stelle:1976gc]. Together, these arguments provide compelling evidence that effective gravitational theories incorporating quantum corrections should generically depend on both $R$ and $P$, with implications for both classical dynamics and BH thermodynamics.
Wald’s entropy formula for a generic action depending on arbitrary curvature invariants reads [DAgostino:2024sgm, DAgostino:2024ymo]
$S_{\text{Wald}}=-2\pi\int_{\mathcal{H}}d^{2}x\sqrt{h}\,\epsilon_{\mu\nu}\epsilon_{\rho\sigma}\frac{\partial\mathcal{L}}{\partial R_{\mu\nu\rho\sigma}}\,$, (24)
where $\epsilon_{\mu\nu}$ is the binormal to the horizon cross-section $\mathcal{H}$, and $\mathcal{L}$ is the Lagrangian density. For the case $\mathcal{L}=f(R,P)$, one has
$\frac{\delta\mathcal{L}}{\delta R_{\mu\nu\rho\sigma}}=\sqrt{-g}\left[f_{R}\frac{\delta R}{\delta R_{\mu\nu\rho\sigma}}+f_{P}\frac{\delta P}{\delta R_{\mu\nu\rho\sigma}}\right],$ (25)
where
$f_{R}\equiv\frac{\partial f}{\partial R},\quad f_{P}\equiv\frac{\partial f}{\partial P}\,.$ (26)
Taking into account the symmetries of the Riemann tensor and the relative contractions,
$R^{\mu}_{\ \alpha\mu\beta}=R_{\alpha\beta}\,,\quad R=g^{\alpha\beta}R_{\alpha\beta}\,$, (27)
one obtains
$\frac{\delta R}{\delta R_{\mu\nu\rho\sigma}}=g^{\mu\rho}g^{\nu\sigma}-g^{\mu\sigma}g^{\nu\rho}\,.$ (28)
Then, the contraction of the binormal tensors with the metric yields [^2] Note the antisymmetry of the binormal, $\epsilon_{\mu\nu}=-\epsilon_{\nu\mu}$, and its normalization on the bifurcation surface, $\epsilon_{\mu\nu}\epsilon^{\mu\nu}=-2$.
$\epsilon_{\mu\nu}\epsilon_{\rho\sigma}\left(g^{\mu\rho}g^{\nu\sigma}-g^{\mu\sigma}g^{\nu\rho}\right)=-2\,.$ (29)
Similarly, for $P=R_{\alpha\beta}R^{\alpha\beta}$, we find
$\frac{\delta P}{\delta R_{\mu\nu\rho\sigma}}=2R^{\alpha\beta}\frac{\delta R_{\alpha\beta}}{\delta R_{\mu\nu\rho\sigma}}\,.$ (30)
Since $R_{\alpha\beta}=R^{\lambda}_{\ \alpha\lambda\beta}$, then
$\frac{\delta R_{\alpha\beta}}{\delta R_{\mu\nu\rho\sigma}}=\frac{1}{2}\left(\delta^{\mu}_{\alpha}\delta^{\rho}_{\beta}g^{\nu\sigma}-\delta^{\mu}_{\alpha}\delta^{\sigma}_{\beta}g^{\nu\rho}+\delta^{\mu}_{\beta}\delta^{\rho}_{\alpha}g^{\nu\sigma}-\delta^{\mu}_{\beta}\delta^{\sigma}_{\alpha}g^{\nu\rho}\right).$ (31)
Hence, one can show that
$\epsilon_{\mu\nu}\epsilon_{\rho\sigma}\frac{\delta R_{\alpha\beta}}{\delta R_{\mu\nu\rho\sigma}}=-2n_{\alpha}n_{\beta}\,$, (32)
where $n^{\mu}$ is the unit normal to $\mathcal{H}$.
Using the above results, the Wald entropy reduces to
$S_{\text{Wald}}=\frac{1}{4G}\int_{\mathcal{H}}d^{2}x\,\sqrt{h}\left[f_{R}+2f_{P}R_{\mu\nu}n^{\mu}n^{\nu}\right].$ (33)
The first term, $f_{R}$, corresponds to the scalar curvature dependence already encountered in $f(R)$ gravity. The second term, proportional to $f_{P}$, encodes the sensitivity of the entropy to local Ricci curvature components normal to the horizon surface. This term can be interpreted as a correction to the gravitational coupling on the horizon, induced by local anisotropies or tidal effects in the geometry.
For static and spherically symmetric BHs, the curvature scalars on the horizon scale as inverse powers of $r_{H}$,
$R\propto\frac{1}{r_{H}^{2}},\quad P\propto\frac{1}{r_{H}^{4}}\,$, (34)
thus, we can write the following relations:
$R=\frac{1}{A}\,,\quad R_{\mu\nu}n^{\mu}n^{\nu}=\frac{\eta}{A}\,$, (35)
where $\eta$ is a dimensionless constant. Notice that the first of the above relations follows from comparison with Eq. (18).
Therefore, Eq. (33) reads
$S_{\text{Wald}}=\frac{A}{4G}f_{R}\left(R,P\right)+\frac{\eta}{2G}f_{P}\left(R,P\right).$ (36)
We thus impose the condition
$\frac{A}{4G}f_{R}(R,P)+\frac{\eta}{2G}f_{P}(R,P)=S_{\text{GUP}}(A)\,$, (37)
which can be written as
$f_{R}+\alpha Rf_{P}=h(R)\,$, (38)
where $\alpha\equiv{2\eta}$ and
$h(R)\equiv{4GR}\hskip 0.85358ptS_{\text{GUP}}\,.$ (39)
Employing Eq. (13) with $A=1/R$, we can see that $h(R)$ coincides with Eq. (21).
The characteristic curves satisfying Eq. (38) are
$\frac{dP}{dR}=\alpha R\,$, (40)
which integrates to
$P=\frac{\alpha}{2}R^{2}+C\,$, (41)
where $C$ is a constant that identifies the particular characteristic curve. Along each such curve, we have
$\frac{df}{dR}=h(R)\,$, (42)
hence, the general solution to Eq. (38) takes the form
$f(R,P)=\int h(R)\,dR+\chi(x)\,$, (43)
where $\chi(x)$ is an arbitrary function of $x\equiv P-\frac{\alpha}{2}R^{2}$. Notice that, for the particular case $\chi(x)=0$, one exactly recovers Eq. (22).
By defining $F(R)\equiv\int h(R)dR$, Eq. (43) reads
$f(R,P)=F(R)+\chi\big(x(R,P)\big).$ (44)
Then, from Eq. (36), we have
$S_{\text{Wald}}=\frac{A}{4G}F^{\prime}+\left(\frac{\eta}{2G}-\frac{A\alpha R}{4G}\right)\chi^{\prime}\,$, (45)
where $F^{\prime}\equiv dF/dR$ and $\chi^{\prime}\equiv d\chi/dx$. Now, we note that
$\frac{\eta}{2G}-\frac{A\alpha R}{4G}=0\,$, (46)
in view of Eqs. (35) and (39). We can conclude that the function $\chi$ does not contribute to the entropy, whose final expression is then given by Eq. (13).
Our findings indicate that the gravitational Lagrangian $f(R,P)$ cannot be entirely determined from Wald’s entropy. Therefore, $\chi(x)$ remains unconstrained by thermodynamic arguments alone, and must be determined via other physical arguments, e.g., the requirement of regularity, stability of solutions, cosmological evolution, or compatibility with quantum corrections. Although the covariant curvature invariant $P$ does not enter the entropy expression, it does contribute to the gravitational field equations and can influence the spacetime dynamics, potentially leading to observable effects that are entirely decoupled from BH thermodynamics.
IV Coefficient Mapping and Physical Interpretation
The correspondence between entropy corrections induced by the GUP and the structure of the modified gravitational Lagrangian allows for a direct interpretation of higher-curvature contributions in terms of underlying quantum gravitational effects.
GUP index $k$ Entropy correction GR correction
1 $\sqrt{A}$ $R^{3/2}$
2 $\ln A$ $R^{2}\left(\ln R^{2}-1\right)$
$\geq 3$ $A^{1-k/2}$ $R^{1+k/2}$
Table 1: Correspondence between GUP modifications and gravity corrections.
We summarize our results in Table 1. As discussed in Sec. II.1, when restricting to leading-order corrections in $c_{k}$, the case $k=1$ yields a contribution proportional to $\sqrt{A}$, corresponding to an $R^{3/2}$ correction in the action, which may indicate IR modifications or non-local effects. The $k=2$ term produces a logarithmic correction to the entropy, associated with the structure $R^{2}(\ln R^{2}-1)$ in the action. Higher-order terms with $k\geq 3$ give rise to progressively suppressed entropy corrections, mapping to higher powers of $R$ ($\sim R^{1+k/2}$), and thus become relevant primarily in UV regimes of large curvature.
When considering the more general Lagrangian $f(R,P)$, we demonstrated that the underlying gravity theory includes an undetermined function $\chi(x)$, with $x=P-\frac{\alpha}{2}R^{2}$, that remains invisible to entropy but can influence spacetime dynamics.
For the sake of concreteness, we restrict the analysis to GUP corrections up to second order, i.e.
$\Xi(\Delta p)=c_{1}\ell_{P}\Delta p+c_{2}(\ell_{P}\Delta p)^{2}\,$, (47)
which leads to
$\Delta p=\frac{1}{2\Delta x}\left(1+\frac{c_{1}\ell_{P}}{2\Delta x}+\frac{\ell_{P}^{2}(c_{1}^{2}+c_{2})}{4\Delta x^{2}}\right).$ (48)
In this case, the entropy reads
$S_{\text{GUP}}=\frac{A}{4G}-\frac{1}{4\zeta}\left(c_{1}\sqrt{\frac{A}{G}}+\frac{c_{1}^{2}+c_{2}}{4\zeta}\ln A\right).$ (49)
Notably, the appearance of both $\sqrt{A}$ and $\ln A$ corrections is in agreement with a wide range of results found in the literature [Carlip:2000nv, Kaul:2000kf, Sen:2012dw, Adler:2001vs, Medved:2004yu, Nozari:2008rc, Majumder:2012rtc, Tawfik:2015kga].
From Eq. (49), we obtain a modified gravity theory of the form
$f(R)=R-\frac{2G}{3\zeta}\left[\frac{c_{1}}{\sqrt{G}}R^{3/2}+\frac{3(c_{1}^{2}+c_{2})}{32\zeta}R^{2}\left(1-\ln R^{2}\right)\right],$ (50)
which introduces higher-order and non-polynomial curvature terms.
The presence of the $R^{3/2}$ and $R^{2}\ln R^{2}$ terms is of particular interest. The former reflects the impact of linear GUP corrections, while the latter combines the effects of both linear and quadratic deformations. Such non-analytic curvature contributions often emerge in quantum corrections to the effective action, e.g., from trace anomalies or conformal field theories on curved backgrounds [Duff:1993wm, Birrell:1982ix, Avramidi:2000bm], as well as from gravity models with entropy-based reconstruction schemes [Cognola:2005de]. This further supports the idea that GUP-inspired corrections may capture essential features of an underlying quantum theory of gravity.
V Astrophysical application
As an astrophysical application of the framework developed above, we now consider an extended gravity theory inspired by the Lagrangian (50), where we restore the radial dependence of the Ricci scalar $R$. For this purpose, by varying the action with respect to the metric tensor, we obtain the $f(R)$ field equations [Sotiriou:2008rp]:
$R_{\mu\nu}f_{R}-\frac{1}{2}g_{\mu\nu}f-(\nabla_{\mu}\nabla_{\nu}{}-g_{\mu\nu}\Box)f_{R}=8\pi GT_{\mu\nu}\,$, (51)
where $\Box\equiv\nabla^{\mu}\nabla_{\mu}$ is the D’Alembert operator and $T_{\mu\nu}$ is the energy-momentun tensor of matter fields. These equations reduce to Einstein’s ones as soon as $f(R)=R$. Taking the trace of Eq. (51), one finds
$Rf_{R}-2f+3\Box f_{R}=8\pi G\,T\,.$ (52)
In general, solving the $f(R)$ field equations may be highly challenging. For the sake of simplicity, we thus consider the quadratic GUP model, i.e., $c_{k}=0$ for $k\neq 2$. This implies
$\Delta x\Delta p=\frac{1}{2}\left(1+c_{2}\ell_{P}^{2}\Delta p^{2}\right).$ (53)
From Eq. (50) (with $c_{1}=0$), the above assumption translates to the modified gravity Lagrangian
$f(R)=R-\frac{Gc_{2}}{16\zeta^{2}}R^{2}\left(1-\ln R^{2}\right)\approx R-\epsilon R^{2}\,$, (54)
where we have defined
$\epsilon\equiv\frac{Gc_{2}}{16\zeta^{2}}\,.$ (55)
In the last step of Eq. (54), we have neglected the $R^{2}\ln R^{2}$ term, which is sub-leading with respect to the $R^{2}$ term at astrophysical scales [^3] We remind that Eq. (54) is up to an integration constant that appears as $\epsilon R^{2}\ln\mu$, where $\mu$ has dimensions of $R^{-2}$. In our context, $\mu$ can be fixed such that $\mu R^{2}\sim 1$ at astrophysical scales.. This framework corresponds to the well-known Starobinsky model [Starobinsky:1980te], which reproduces GR in the limit where the GUP reduces to the Heisenberg principle, namely for $c_{2}\rightarrow 0$.
We can attempt to solve the resulting differential problem in the weak-field slow-motion limit of the theory [Schmidt:1986via, Capozziello:2009zz, Berry:2011pb]. To do that, we search for solutions under the following metric ansatz:
$ds^{2}=(1-2\phi)dt^{2}-(1+2\psi)(dr^{2}+r^{2}d\Omega^{2})\,$, (56)
where $d\Omega^{2}\equiv d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}$. In this approach, we can neglect time derivatives and products of the potentials $\phi$ and $\psi$. Then, one finds
$R=2(2\nabla^{2}\psi-\nabla^{2}\phi)\,$, (57)
where $\nabla^{2}\equiv\delta^{ij}\partial_{i}\partial_{j}$. In addition, we have $T_{\mu\nu}=\delta_{\mu}^{0}\delta_{\nu}^{0}\rho$, where $\rho$ is the rest-mass energy density. Hence, from the (00)-component of Eq. (51) and from Eq. (52), one finds
$-\nabla^{2}\phi-\frac{R}{2}+4\epsilon\nabla^{2}R$ $=8\pi G\,$, (58)
$-R+12\epsilon\nabla^{2}R$ $=8\pi G\,$, (59)
respectively. Combining Eqs. (58) and (59), we obtain
$\nabla^{2}\phi+2\epsilon\nabla^{2}R=-4\pi G\rho\,.$ (60)
After some manipulations, we arrive at
$(1-12\epsilon\nabla^{2})\nabla^{2}\phi=-4\pi G(1-16\epsilon\nabla^{2})\rho\,.$ (61)
The above equation represents a fourth-order gravity theory motivated by quantum considerations. Therefore, the solutions must guarantee a Newtonian behavior in the long-range regime, i.e., they should be exponentially vanishing [Stelle:1977ry]. In particular, taking $\rho$ as a delta source in Eq. (61), i.e. $\rho=M\delta$, we get the Yukawa-like solutions
$\phi(r)$ $=\frac{GM}{r}\left(1+\frac{e^{-r/\sqrt{12\epsilon}}}{3}\right),$ (62)
$\psi(r)$ $=\frac{GM}{r}\left(1-\frac{e^{-r/\sqrt{12\epsilon}}}{3}\right),$ (63)
with $\epsilon>0$ to avoid oscillating behaviors. In light of Eq. (55), this condition implies $c_{2}>0$. From Eq. (53), this is precisely the requirement for the existence of a minimal length [Kempf:1994su]. It is straightforward to verify that Eqs. (62)–(63) reproduce the linearized Schwarzschild metric in the regime $r\gg\sqrt{12\epsilon}$, i.e., at distances well above the Planck scale.
We can now examine the observable consequences of the GUP-corrected metric by analyzing the light deflection by a massive body. To this aim, we follow the treatment of [Weinberg:1972kfs, Scardigli:2014qka] and consider a photon approaching the Sun from very large distances. Then, the photon orbit $\vartheta(r)$ in given by
$\vartheta(r)=\int_{\infty}^{r}dr^{\prime}\,\mathcal{I}(r^{\prime})\,$, (64)
where
$\mathcal{I}(r)\equiv\frac{1}{r}\left[\left(\frac{r}{r_{0}}\right)^{2}\mathcal{F}(r_{0})-\mathcal{F}(r)\right]^{-1/2},$ (65)
with $\mathcal{F}(r)=1-2\phi(r)$, and $r_{0}$ being the minimum distance between the Sun and the photon. The total deflection angle is then calculated as
$\Delta\vartheta=2|\gamma(r_{0})|-\pi\,.$ (66)
To simplify calculations, we write
$\mathcal{I}(r)=\frac{1}{r}\left[\left(\frac{r}{r_{0}}\right)^{2}-1\right]^{-1/2}\left[\mathcal{F}(r_{0})+\frac{\mathcal{F}(r_{0})-\mathcal{F}(r)}{(r/r_{0})^{2}-1}\right]^{-1/2}.$ (67)
We can write the argument of the last squared bracket as
$\mathcal{F}(r_{0})+\frac{\mathcal{F}(r_{0})-\mathcal{F}(r)}{(r/r_{0})^{2}-1}=1-\delta\,$,
where the contribution
$\delta\equiv\frac{2GM\left[r^{3}\left(3+e^{-r_{0}/\sqrt{12\epsilon}}\right)-r_{0}^{3}\left(3+e^{-r/\sqrt{12\epsilon}}\right)\right]}{3rr_{0}\left(r^{2}-r_{0}^{2}\right)}$ (68)
is expected to be subdominant compared to unity in the weak field regime, i.e. $GM/r_{0}\ll 1$. We can thus employ the expansion $(1-\delta)^{-1/2}\simeq 1+\delta/2+3\delta^{2}/8$ and recast Eq. (67) as
$\mathcal{I}(r)\simeq\mathcal{I}_{1}(r)+\mathcal{I}_{2}(r)+\mathcal{I}_{3}(r)\,$, (69)
where
$\mathcal{I}_{1}(r)\equiv\frac{r_{0}}{r\sqrt{r^{2}-r_{0}^{2}}}\,$, (70a)
$\mathcal{I}_{2}(r)\equiv\frac{GM\left[r^{3}\left(3+e^{-r_{0}/\sqrt{12\epsilon}}\right)-r_{0}^{3}\left(3+e^{-r/\sqrt{12\epsilon}}\right)\right]}{3r^{2}\left(r^{2}-r_{0}^{2}\right)^{3/2}}\,$, (70b)
$\mathcal{I}_{3}(r)\equiv\frac{G^{2}M^{2}\left[r^{3}\left(3+e^{-r_{0}/\sqrt{12\epsilon}}\right)-r_{0}^{3}\left(3+e^{-r/\sqrt{12\epsilon}}\right)\right]^{2}}{6r^{3}r_{0}\left(r^{2}-r_{0}^{2}\right)^{5/2}}.$ (70c)
Then, we have
$\int_{\infty}^{r_{0}}dr\,\mathcal{I}_{1}(r)=-\frac{\pi}{2}\,$, (71)
while the integrals $\mathcal{I}_{2}(r)$ and $\mathcal{I}_{3}(r)$ are not as straightforward. Since $\sqrt{\epsilon}/r\ll 1$ throughout the entire range of integration (we recall that $\sqrt{\epsilon}\propto\ell_{P}$), it is evident that, for fixed $\epsilon$, the exponential factors $e^{-r/\sqrt{12\epsilon}}$ in Eqs. (70b) and (70c) become increasingly subdominant as $r$ grows relative to $r_{0}$. In other words, the only non-negligible contributions of these exponentials to the integral (64) are expected to come from radii close to the minimal value $r_{0}$. Hence, in Eqs. (70b) and (70c), we may approximate $e^{-r/\sqrt{12\epsilon}}\approx e^{-r_{0}/\sqrt{12\epsilon}}$, which ensures that the limit for $r\to r_{0}^{+}$ is finite [^4] The divergences as $r\to r_{0}^{+}$ cancel out due to the opposite signs of the terms in Eqs. (70b) and (70c).. In this case, we find
$\int_{\infty}^{r_{0}}dr\,\mathcal{I}_{2}(r)\approx-\frac{GM}{r_{0}}\left(1+\frac{e^{-r_{0}/\sqrt{12\epsilon}}}{3}\right),$ (72)
$\int_{\infty}^{r_{0}}dr\,\mathcal{I}_{3}(r)\approx-\left(\frac{GM}{r_{0}}\right)^{2}\left(1+\frac{e^{-r_{0}/\sqrt{12\epsilon}}}{3}\right)^{2}\left(\frac{15\pi}{8}-2\right).$ (73)
Therefore, from Eq. (66), we finally obtain
$\Delta\vartheta\approx$ $\ \frac{4GM}{r_{0}}\left(1+\frac{e^{-r_{0}/\sqrt{12\epsilon}}}{3}\right)+$
$\left(\frac{4GM}{r_{0}}\right)^{2}\left(\frac{15\pi}{64}-\frac{1}{4}\right)\left(1+\frac{e^{-r_{0}/\sqrt{12\epsilon}}}{3}\right)^{2}.$ (74)
It easy to verify that, in the limit $\epsilon\to 0^{+}$, Eq. (74) reduces to
$\Delta\vartheta\approx\frac{4GM}{r_{0}}+\left(\frac{GM}{r_{0}}\right)^{2}\left(\frac{15\pi}{4}-\frac{1}{4}\right),$ (75)
which coincides with the prediction of the Schwarzschild metric arising from the second-order expansion in $GM/r_{0}\ll 1$  [^5] In the literature, Eq. (75) is often expressed as [Epstein:1980dw, Fischbach:1980su, Richter:1982zz] $\Delta\vartheta\approx\frac{4GM}{b}+\frac{15\pi}{4}\left(\frac{GM}{b}\right)^{2},$ (76) where $b=r_{0}/\sqrt{1-2GM/r_{0}}$ is the impact parameter. One can readily check that Eqs. (75) and (76) agree up to second order in $GM/r_{0}\ll 1$..
It is now interesting to compare our theoretical result with the measurements of the deflection of light by the Sun. In the case of a photon just grazing the Sun’s surface, one usually has
$\Delta\vartheta_{\odot}=\frac{4GM_{\odot}}{R_{\odot}}\left(\frac{1+\gamma}{2}\right),$ (77)
where $GM_{\odot}\approx 7.5\times 10^{9}\ \text{eV}^{-1}$ and $R_{\odot}\approx 3.5\times 10^{15}\,\text{eV}^{-1}$. Here, $\gamma$ is the standard post-Newtonian parameter, whose value has been measured, e.g., with the very long baseline interferometry (VLBI) [Lambert:2009xy]:
$|\gamma-1|\lesssim 1.6\times 10^{-4}\,.$ (78)
Substituting Eq. (78) into Eq. (77) and comparing with Eq. (74), we obtain
$\epsilon\lesssim 1.5\times 10^{28}\,\text{eV}^{-2}\,.$ (79)
Finally, a rough estimate of the quadratic GUP correction (53) follows from Eq. (55), from which we get $c_{2}=16\epsilon\zeta^{2}G^{-1}$. Recalling that $\zeta=\Delta x\,A^{-1/2}$ (see below Eq. (8)) and assuming $\Delta x\sim\ell_{P}$, we find $c_{2}\sim 16\epsilon A^{-1}$. Since in our framework $A$ must be interpreted as the horizon area of a solar-mass BH, i.e., $A=4\pi R_{S,\odot}^{2}$ with $R_{S,\odot}\approx 1.5\times 10^{10}\,\text{eV}^{-1}$, we finally obtain
$c_{2}\sim\frac{4\epsilon}{\pi R_{S,\odot}^{2}}\lesssim 8.5\times 10^{7}\,.$ (80)
From Eq. (53), we know that $\Delta x_{\text{min}}=\sqrt{c_{2}}\ell_{P}$ [Kempf:1994su], we get $\Delta x_{\text{min}}\lesssim 9.2\times 10^{3}\ell_{P}$. It is worth emphasizing that the constraint (80) provides a stringent limit on the quadratic GUP model, improving upon previous results in the literature (see, e.g., [Bosso:2023aht] and references therein).
VI Conclusions
In this work, we have developed a comprehensive framework that establishes an explicit connection between GUP-inspired quantum corrections and extended gravitational actions containing higher-curvature terms. Starting from a generic deformation of the Heisenberg uncertainty principle, expressed as a power series in the momentum uncertainty $\Delta p$, we showed that the resulting quantum corrections manifest as modifications to BH entropy. Using Wald’s Noether-charge formalism, we systematically mapped these entropy corrections onto corresponding deformations of the gravitational Lagrangian within the $f(R)$ theories. We further extended the analysis to the $f(R,R_{\mu\nu}R^{\mu\nu})$ framework, motivated by the fact that quantum field theory in curved spacetime and effective action computations naturally generate a variety of curvature invariants beyond the Ricci scalar alone. Consequently, the resulting effective Lagrangians capture a more realistic and nuanced picture of gravitational dynamics near Planckian regimes.
Our theoretical construction reveals a detailed correspondence between the coefficients characterizing the GUP deformation and those appearing in the BH entropy corrections, ultimately linking them to the parameters governing the effective gravitational action. This hierarchical mapping bridges the gap between phenomenological quantum gravity models, which often introduce generalized uncertainty principles heuristically, and semi-classical modified gravity theories that encode quantum corrections at the level of the classical action.
We also examined potential astrophysical implications of the obtained GUP-inspired modified gravity models. In particular, we showed that the second-order GUP model leads to solutions characterized by Yukawa-like corrections to the standard Schwarzschild metric in the weak field regime. Then, we tested our theoretical framework by analyzing light deflection by the Sun. By comparing the total deflection angle, expanded to second order in the weak-field regime, with current experimental bounds, we obtained a very stringent constraint on the quadratic GUP model.
Looking ahead, several directions naturally extend this research. Investigating BH solutions within our framework and comparing predictions with observational data, such as gravitational wave signals or BH imaging, would provide concrete tests of the theoretical constructs developed here. Furthermore, exploring the cosmological implications of the obtained GUP-inspired modified gravity models could shed light on their potential role in early universe phenomena, inflationary dynamics, or the present accelerated expansion. Such studies may open new observational or experimental avenues for probing aspects of quantum gravity phenomenology, particularly in models featuring a minimal length scale.
Acknowledgments
R.D. acknowledges financial support from INFN-Sezione di Roma 1, esperimento Euclid. The research of G.G.L. is supported by the Ayuda Postdoctoral of the University of Lleida (project code: X25027). G.G.L. gratefully acknowledges the contribution from Cost Action CA21106 – COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers). P.B. and G.G.L. acknowledge the contribution of the LISA Cosmology Working Group (CosWG), as well as support from Cost Action CA23130 – Bridging high and low energies in search of quantum gravity (BridgeQG). R.D. and G.G.L. acknowledge work from COST Action CA21136 – Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by European Cooperation in Science and Technology.
Appendix A $\Delta p$ series inversion
In this appendix, we aim to invert Eq. (3) in order to find an expansion that would describe $\Delta p$ in terms of powers of $\Delta x$. Let us thus rewrite Eq. (3) as
$\Delta p=\sum_{n=0}^{\infty}\ell_{P}^{n}\Delta p_{n}\,.$ (81)
This equation can be read as a power series in $\ell_{P}$. Equating the coefficients of a specific power $n$ of $\ell_{P}$ from the left and right hand sides of Eq. (3), using Eq. (81), we can write
$\Delta p_{n}=\frac{1}{2\Delta x}\sum_{k=1}^{n}c_{k}\sum_{\vec{\mu}\in P^{(k)}(n-k)}\binom{k}{\vec{\mu}}\prod_{j=0}^{n-k}\Delta p_{j}^{\mu_{j}}\,$, (82)
where $P^{(k)}(n-k)$ is the set of the integer partitions of nkn-k in $k$ parts, that is the set of $k$ non-negative integers whose sum gives nkn-k, while we use the notation $\vec{\mu}$ to designate the set of integers $\mu_{j}\in\{0,n-k\}$, corresponding to the multiplicities for such partitions, that is $n-k=0\,\mu_{0}+1\,\mu_{1}+\dots+n\,\mu_{n}$ and $\sum_{j=0}^{n}\mu_{j}=k$. Furthermore, the symbol $\binom{k}{\vec{\mu}}$ stands for the following multinomial
$\binom{k}{\vec{\mu}}=\binom{k}{\mu_{0},\mu_{1},\dots,\mu_{n}}=\frac{k!}{\mu_{0}!\mu_{1}!\dots\mu_{n}!}\,.$ (83)
It is possible to invert the relation above, obtaining
$\Delta p_{n}=\frac{1}{(2\Delta x)^{n+1}}\sum_{\vec{m}\in P(n)}\frac{n^{\underline{|\vec{m}|-1}}}{\vec{m}!}\vec{c}^{\vec{m}},$ (84)
where $P(n)$ is the set of integer partitions of $n$, corresponding to the set of any length of positive-definite integers whose sum gives $n$. Here, we have adopted the following notation:
  1. (i)
    $\vec{m}\in P(n)$ is the array of positive-definite integers such that $\sum_{j}j\,m_{j}=n$;
  2. (ii)
    $|\vec{m}|=\sum_{j}m_{j}$;
  3. (iii)
    $\vec{m}!=\prod_{j}m_{j}!$;
  4. (iv)
    $\vec{c}^{\vec{m}}=\prod_{j}c_{j}^{m_{j}}$;
  5. (v)
    $n^{\underline{m}}=n(n-1)\dots(n-m+1)$ is a falling factorial.
In order to prove Eq. (84), let us introduce the following generating functions:
$\Delta P(t)$ $=\sum_{n=0}^{\infty}\Delta p_{n}t^{n},$ $C(t)$ $=\sum_{n=0}^{\infty}c_{n}t^{n},$ (85)
with $\Delta p_{0}=1/(2\Delta x)$ and $c_{0}=1$. Using the expansion as in Eq. (82), we can show that
$\Delta P(t)=\frac{1}{2\Delta x}\left[1+C\left(t\Delta P(t)\right)\right].$ (86)
At this point, it is convenient to introduce
$Y(t)=t\Delta P(t)\,$, (87)
so that the relation above can be written as
$t=\frac{2\Delta x\,Y}{1+C(Y)}\,.$ (88)
Furthermore, we notice that
$Y(t)=\frac{1}{2\Delta x}+\sum_{n=0}\Delta p_{n}t^{n+1}\,.$ (89)
Therefore, inverting Eq. (88) in light of Eq. (89) would give us the expression for $\Delta p_{n}$. In order to achieve this goal, we employ the Lagrange inversion theorem, which in our case can be spelled as
$\Delta p_{n}$ $=[t^{n+1}]Y(t)$
$=\frac{1}{(n+1)(2\Delta x)^{n+1}}[Y^{n}]\left(1+C(Y)\right)^{n+1},$ (90)
where we use the notation $x^{n}]f(x)$ to denote the coefficient of $x^{n}$ in $f(x)$. Thus, we observe that
$(1+C(Y))^{n+1}=\sum_{N=0}^{\infty}\sum_{\vec{\nu}\in P^{(n+1)}(N)}\binom{n+1}{\vec{\nu}}Y^{N}\vec{c}^{\vec{\nu}}.$ (91)
At this point, since $|\vec{\nu}|=n+1$ and given that the tuple $\vec{\nu}$ may also contain an element $\nu_{0}$ corresponding to the multiplicity of 0 in the partition of $N$, we have
$\binom{n+1}{\vec{\nu}}=\frac{(n+1)^{\underline{|\vec{\nu}|}}}{\vec{\nu}!}=(n+1)\frac{n^{\underline{\nu_{1}+\nu_{2}+\dots-1}}}{\nu_{1}!\nu_{2}!\dots}\,.$ (92)
Hence, we conclude
$\Delta p_{n}$ $=\frac{1}{(2\Delta x)^{n+1}}\sum_{\vec{\nu}\in P^{(n+1)}(n)}\frac{n^{\underline{\nu_{1}+\nu_{2}+\dots-1}}}{\nu_{1}!\nu_{2}!\dots}\vec{c}^{\vec{\nu}}$
$=\frac{1}{(2\Delta x)^{n+1}}\sum_{\vec{m}\in P(n)}\frac{n^{\underline{|\vec{m}|-1}}}{\vec{m}!}\vec{c}^{\vec{m}}\,.$ (93)